OpenStudy (anonymous):

rearange the equations.... b+4=a b= 4d=c d= m-3=4k m=

OpenStudy (anonymous):

1. subtract 4 from each side b = a-4 2. divide both sides by 4 d = c/4 3. add 3 to both sides m = 4k+3

OpenStudy (anonymous):

b=a-4 d=c/4 m=4k+3

OpenStudy (anonymous):

well said juanita an fontez